OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3? y = 3x + 5 y = 3x + 3 y = 3x - 17 y = 3x - 5

hero (hero):

You do realize that (x,y) = (1,6) right?

OpenStudy (anonymous):

yes

hero (hero):

So if you know that, all you need to do is find the right b value. What prevents you from setting up your own equation of the form y = mx + b and include all of the given info? In this case, x = 1, y = 6, and m = 3. Input those given values into y = mx + b and then solve for b

OpenStudy (anonymous):

or use the formula (y-y2)=m(x-x2) " hello hero" :D

OpenStudy (anonymous):

ok

hero (hero):

All he needs to do is find "b" @magepker728. He doesn't need to do more work than is necessary.

hero (hero):

Oh, and "Hi" back at you @magepker728

OpenStudy (anonymous):

so it's either the first one or the last one

OpenStudy (anonymous):

right